solve using quadratic formula:
1/4x^2+17/4=2x

Answer by stanbon(75887)

solve using quadratic formula:
1/4x^2+17/4=2x
---
Multiply thru by 4 to get rid of the fractions:
x^2 + 17 = 8x
---
Rearrange:
x^2 - 8x + 17 = 0
Use Quadratic Formula:
---
x = [8 +- sqrt(64 -4*17]/2
---
x = [8 +- sqrt(-4)]/2
---
x = [8 +- 2i]/2
---
x = [4 + i] or x = [4-i]
============================
Cheers,
Stan H.